OpenStudy (anonymous):

Use the method of substitution to solve the system. (If there is no solution, enter NO SOLUTION.) x^2 + z^2 = 17 4x + y = 5 y + z = 5

OpenStudy (aum):

x^2 + z^2 = 17 ----- (1) 4x + y = 5 ----- (2) y + z = 5 ----- (3) From (3): y = 5 - z Put this in (2): 4x + 5 - z = 5 subtract 5: 4x - z = 0 add z: 4x = z or z = 4x. Put this in (1): x^2 + (4x)^2 = 17 solve for x Then solve for z and y.